# On the stability of global forecasting models

### v1

@@ Abstract
Forecast stability, that is the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems. Despite its importance, stability is often overlooked in favor of accuracy, particularly in global forecasting models. In this study, we evaluate the stability of point and probabilistic forecasts across different retraining frequencies and ensemble strategies using two large retail datasets (M5 and VN1). To do this, we introduce a new metric for probabilistic stability (MQC) and analyze ten different global models and four ensemble configurations. The results show that less frequent retraining not only preserves but often improves forecast stability, while ensembles, especially those combining diverse pool of models, further enhance consistency without sacrificing accuracy. These findings challenge the need for continuous retraining and highlight ensemble diversity as a key factor in reducing forecast stability. The study promotes a shift toward stability-aware forecasting practices, offering practical guidelines for building more robust and sustainable prediction systems.
@@ 1 Introduction
In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand forecasting. While their accuracy and efficiency have been widely studied, less attention has been paid to the stability of their forecasts over time. In many operational contexts, forecasts are produced regularly and serve as the foundation for critical business decisions, from inventory planning to resource allocation. In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably consistent over time. Indeed, forecast stability refers to the consistency of predictions produced by a forecasting model as new data becomes available. As defined by Godahewa et al. (2025), forecast stability can be categorized into two primary forms: vertical stability, which concerns the consistency of forecasts for the same target date across different forecast origins, and horizontal stability, which addresses the smoothness of forecasts across the forecast horizon from a single origin. Both types play a distinct role: vertical stability helps avoid costly forecast revisions that can disrupt planning cycles, while horizontal stability prevents erratic behavior across time steps that could lead to inefficient operations and amplification of demand fluctuations, such as the bullwhip effect in supply chains (Lee et al., 1997). In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is robust to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions. Forecast instability, indeed, can have serious consequences. Unstable forecasts may lead to frequent and expensive adjustments in supply chain plans, diminish trust in the forecasting system, and complicating the decision-making processes, leading to suboptimal business outcomes.
@@
Nevertheless, despite the importance of forecast stability, it is still common practice to evaluate forecasting models only on their accuracy, even in frequent retraining settings where stability is more likely to be compromised. This can also be due to an inherent perception that there exists a trade-off between accuracy and stability. As new information becomes available, updated forecasts may naturally differ from earlier versions, ideally improving in accuracy at the expense of stability. Conversely, overly stable forecasts might sacrifice accuracy by ignoring valuable new information. Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance while enhancing stability (Wang et al., 2023). By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability. Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational consistency, recent research suggests that stability and accuracy are not necessarily in conflict, and both can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)). Moreover, Zanotti (2025b) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, meaning that models estimated using this forecasting approach retain the same level of performance even when updated less frequently. This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because stability should be a non-increasing function of the retraining frequency.
@@ 1.1 Research Question
We aim to address the question ”Are global forecasting models stable?”. Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is whether avoiding re-estimation for every new observation damages the stability of global models. To address this question, we rely on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition data.
@@ 1.2 Contributions
    We analyze the relationship between retraining frequency and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
@@
    We suggest a new metric to evaluate the instability of probabilistic forecasts.
@@ 2 Related works
The cross-learning approach has seen substantial development in recent years. Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field. Semenoglou et al. (2021) demonstrated the high accuracy of GMs on the M4 competition dataset, while Hewamalage et al. (2022) explored the conditions under which global models are competitive. Additionally, Montero-Manso & Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of local models, with lower complexity and without assuming data similarity. Global models have proven to be the most accurate method in several forecasting domains, including retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), gas consumption (Gaweł & Paliński, 2024), and crop production (Ibañez & Monterola, 2023). Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022a), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022). To further enhance GM performance, several strategies such as clustering (Godahewa et al. (2021a), Bandara et al. (2020)) and data augmentation (Bandara et al., 2021) have been explored. Moreover, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures have been specifically designed to support cross-learning. Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their interpretability (Rajapaksha et al., 2023).
@@
From a forecasting stability perspective, the literature is very poor. Godahewa et al. (2025) introduced the first classification of the different types of forecasting stability, that is horizontal or vertical stability, proposing a new model-agnostic framework based on linear combinations of predictions to obtain more stable forecasts. Van Belle et al. (2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective as well as a forecast stability perspective, directly including an instability component into the loss function of the model. There is also active research on forecast stability within judgmental forecasting ground (Fildes & Goodwin, 2021). However, almost all the literature related to forecast stability is focused on point prediction stability only, possibly because it is not clear how to measure instability in terms of probabilistic forecasting, and maybe also because most machine learning and deep learning methods do not directly output probabilistic predictions (Makridakis et al., 2022c). Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). Among the methods developed for uncertainty quantification, Vovk et al. (2005) introduced Conformal Inference, a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
@@
In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing the retraining frequency of global models on both accuracy and forecasting costs. Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance. However, their focus was limited to the exponential smoothing family, following the traditional local modeling approach. (Huber & Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset. While these findings are encouraging, there has been little direct investigation into whether global models specifically can retain stability with less frequent updates.
@@ 3.1 Datasets
For our experiments, we employed two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition, part of the well-known M-competitions series organized by Spyros Makridakis and colleagues, aimed to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) includes 3,049 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin. The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state). The dataset also includes exogenous variables such as prices, promotions, and special events (e.g., holidays), which can influence demand. The VN1 Forecasting - Accuracy Challenge, organized in October 2024 by Flieber, Syrup Tech, and SupChains, represents the first edition of a new competition series (Vandeput, 2024). The dataset comprises weekly sales data for 15,053 products sold by U.S.-based e-vendors between 2020 and 2024. Unlike the M5, which involves a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset captures sales from 328 warehouses operated by 46 different retailers. As far as we are aware, our study is among the first to evaluate forecasting models on this dataset. Together, these two datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings.
Table 1: The M5 and the VN1 datasets used in the experiments.
Dataset Frequency N. Series Min Obs per Series
M5 Daily (7) 28.298 730
VN1 Weekly (52) 15.053 157
In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of reducing retraining frequency are likely to be most pronounced. To ensure consistency with the evaluation setup described in Section 3.3, we applied a filtering criterion: for the daily data (M5), only time series with at least two years of observations (730 data points) were retained; for the weekly data (VN1), we considered only those series with at least three years of data (157 observations).
@@ 3.2 Forecasting models
Following Zanotti (2025b), let us define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series. Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model. Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$. Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series. This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
@@
In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining. Cross-learning has become the standard in many industries dealing with large-scale time series data, particularly in retail demand forecasting, where regular forecasts for thousands of SKUs are required (Januschowski et al., 2020). Therefore, to conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures. The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
@@
As in Zanotti (2025b), we used five machine learning models and five deep learning models. Machine learning models have proven effective in forecasting tasks due to their ability to capture complex non-linear relationships in data. They are also relatively easy to train and tend to deliver strong performance when leveraging cross-learning techniques. In this study, we experimented with Linear (Pooled) Regression and four widely used tree-based methods. While machine learning models are easier to train compared to deep learning alternatives, they often rely heavily on high-quality feature engineering (Januschowski et al., 2022). For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M5 and VN1 solutions. We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week). We also included static metadata such as store, product, category, and location identifiers based on the dataset’s frequency. For the M5 dataset, additional external features like special events were also incorporated. Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective libraries. We compared five machine learning models for time series forecasting. Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso & Hyndman, 2021), (Godahewa et al., 2021b)). Random Forest (RF), an ensemble learning method, captures non-linear patterns and is robust to overfitting, making it suitable for demand forecasting ((Breiman, 2001), (Januschowski et al., 2022)). Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen & Guestrin, 2016), (Ke et al., 2017), Makridakis et al. (2022a)). Categorical Boosting (CatBoost) specializes in handling categorical data with minimal preprocessing (Prokhorenkova et al., 2018).
@@
Deep learning models have gained significant traction in time series forecasting due to their ability to model long-range temporal dependencies and to learn hierarchical representations directly from raw input data (Goodfellow et al., 2016). Unlike machine learning models, deep learning architectures typically do not require extensive manual feature engineering. They can autonomously learn lagged structures, trend, and seasonality directly from the raw series. Nevertheless, training deep models is often more complex due to the higher number of hyperparameters and their sensitivity to configuration choices, which can significantly affect performance (Smyl, 2020). In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates as inputs, while using top solutions’ guidelines for setting the model hyperparameters. We compared five deep learning models for time series forecasting. Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter & Schmidhuber, 1997)). Temporal Convolutional Networks (TCN) offer an alternative by using causal convolutions to capture long-range dependencies (Van den Oord et al., 2016). Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
@@
Moreover, we also estimated four different ensemble learning models based on Zanotti (2025a). Ensemble learning combines the predictions of multiple base models to improve forecast accuracy and robustness, particularly when individual models capture complementary patterns in the data. In the context of global forecasting, ensembles are especially valuable for mitigating the risk of model instability and overfitting, which can occur when relying on a single method (Wang et al., 2023). In our study, we adopted a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016). Specifically, we constructed ensembles by averaging the forecasts of the top two, top three, top four, and top five most accurate individual models, as determined on out-of-sample performance. The ensemble configurations were designed to reflect practical setups and were validated using the same rolling origin evaluation framework described in Section 3.3. By incorporating ensemble learning into our analysis, we aimed to assess whether combining global models could further enhance forecasting stability under different retraining scenarios. Moreover, by evaluating multiple ensemble sizes, we aimed to understand how the inclusion of additional models affects the stability under different circumstances.

### v2

@@ Abstract
Forecast stability, that is, the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems. Despite its importance, stability is often overlooked in favor of accuracy. In this study, we evaluate the stability of point and probabilistic forecasts across several retraining scenarios using two large retail demand datasets (M5 and VN1) and ten different global forecasting models. To analyze stability in the probabilistic setting, we propose a new model-agnostic, distribution-free, and scale-free metric that measures probabilistic instability: the Scaled Multi-Quantile Change (SMQC). Furthermore, we also evaluate the effects of retraining on various ensemble configurations based on forecast pooling. The results show that, compared to continuous retraining, less frequent retraining not only preserves but often improves forecast stability, challenging the need for continuous retraining. The study promotes a shift toward stability-aware forecasting practices, proposing a new tool to effectively evaluate forecast stability in probabilistic settings, and offering practical guidelines for building more robust prediction systems.
@@ 1 Introduction
In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand forecasting. While their accuracy and efficiency have been widely studied, less attention has been paid to the stability of their forecasts over time. In many operational contexts, forecasts are produced regularly and serve as the foundation for critical business decisions, from inventory planning to resource allocation. In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably stable over time. Indeed, forecast stability refers to the consistency of predictions produced by a forecasting model as new data becomes available. As defined by Godahewa et al. (2025), forecast stability can be categorized into two primary forms: vertical stability, which concerns the consistency of forecasts for the same target date across different forecast origins, and horizontal stability, which addresses the smoothness of forecasts across the forecast horizon from a single origin. Both types play a distinct role: vertical stability helps avoid costly forecast revisions that can disrupt planning cycles, while horizontal stability prevents erratic behavior across time steps that could lead to inefficient operations and amplification of demand fluctuations, such as the bullwhip effect in supply chains (Lee et al., 1997). In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is stable to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions. Forecast instability, indeed, can have serious consequences. Unstable forecasts may lead to frequent and expensive adjustments in supply chain plans, diminish trust in the forecasting system, and complicating the decision-making processes, leading to suboptimal business outcomes.
@@
Nevertheless, despite the importance of forecast stability, it is still common practice to evaluate forecasting models only on their accuracy, even in frequent retraining settings where stability is more likely to be compromised. This can also be due to an inherent perception that there exists a trade-off between accuracy and stability. As new information becomes available, updated forecasts may naturally differ from earlier versions, ideally improving in accuracy at the expense of stability. Conversely, overly stable forecasts might sacrifice accuracy by ignoring valuable new information. Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance (Wang et al., 2023). By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability. Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational stability, recent research suggests that high levels of stability and accuracy can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)). Moreover, Zanotti (2025b) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, meaning that models estimated using this forecasting approach retain the same level of performance even when updated less frequently. This has a strong potential implication on the forecast stability: if we can reduce the retraining frequency of global models, then we should obtain more stable predictions, because retraining leads to different model artifacts potentially producing stability disruptions.
@@ 1.1 Research Question
We aim to address the question ”Do global models for retail demand forecasting produce stable forecasts?”. Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is whether avoiding re-estimation for every new observation damages the stability of global models. To address this question, we rely on the two most recent and comprehensive retail forecasting datasets: the M5 and VN1 competition data.
@@ 1.2 Contributions
    We analyze the relationship between the retraining period and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
@@
    We suggest new metrics to evaluate the instability of probabilistic forecasts.
@@ 2 Related works
The cross-learning approach has seen substantial development in recent years. Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field.
Following Zanotti (2025b), we define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series. Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model. Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$. Under the local modeling paradigm, forecasts for a horizon $h$ are generated by training a separate model for each individual time series. This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
@@
Semenoglou et al. (2021) demonstrated the high accuracy of GMs on the M4 competition dataset, while Hewamalage et al. (2022) explored the conditions under which global models are competitive. Additionally, Montero-Manso & Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of local models, with lower complexity and without assuming data similarity. Global models have proven to be the most accurate method in several forecasting domains, including retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero & Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß & Hans, 2024), gas consumption (Gaweł & Paliński, 2024), and crop production (Ibañez & Monterola, 2023). Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022a), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022). To further enhance GM performance, several strategies such as clustering (Godahewa et al. (2021a), Bandara et al. (2020)) and data augmentation (Bandara et al., 2021) have been explored. Moreover, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures have been specifically designed to support cross-learning. Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their interpretability (Rajapaksha et al., 2023).
@@
From a forecasting stability perspective, the literature is impoverished. Godahewa et al. (2025) introduced the first classification of the different types of forecasting stability, that is horizontal or vertical stability, proposing a new model-agnostic framework based on linear combinations of predictions to obtain more stable forecasts. Van Belle et al. (2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective, as well as a forecast stability perspective, directly including an instability component into the loss function of the model. There is also active research on forecast stability within judgmental forecasting ground (Fildes & Goodwin, 2021). However, almost all the literature related to forecast stability is focused on point prediction stability only, possibly because it is not clear how to measure instability in terms of probabilistic forecasting, and maybe also because most machine learning and deep learning methods do not directly output probabilistic predictions (Makridakis et al., 2022c). Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). The seminal work of Van Belle et al. (2024) is the only one addressing the problem of probabilistic stability evaluation, but it only applies to a specific forecasting model (NBEATS), which has been modified to output a full predictive Gaussian distribution. Therefore, the stability evaluation of probabilistic forecasts is an open question, especially in a model-agnotic way. Lukily, among the methods developed for uncertainty quantification, Vovk et al. (2005) introduced Conformal Inference, which is a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
@@
In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025b), who showed the positive effects of reducing retraining of global models on both accuracy and forecasting costs. The seminal work of Spiliotis & Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance, focusing on the exponential smoothing family of models, following the traditional local modeling approach. (Huber & Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset. While these findings are encouraging, there has been little direct investigation into whether global models specifically can retain stability with less frequent updates.
@@ 3.1 Datasets
For our experiments, we employed two retail forecasting datasets: the M5 and the VN1 competition datasets. The M5 competition, part of the well-known M-competitions series organized by Spyros Makridakis and colleagues, aimed to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022b). The M5 dataset (Howard & Makridakis, 2020) includes 30,490 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin. The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state). The dataset also includes exogenous variables such as prices, promotions, and special events (e.g., holidays), which can influence demand. The VN1 Forecasting - Accuracy Challenge, organized in October 2024 by Flieber, Syrup Tech, and SupChains, represents the first edition of a new competition series (Vandeput, 2024). The dataset comprises weekly sales data for 15,053 products sold by U.S.-based e-vendors between 2020 and 2024. Unlike the M5, which involves a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset captures sales from 328 warehouses operated by 46 different retailers. As far as we are aware, our study is among the first to evaluate forecasting models on this dataset. Together, these two datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings.
In both cases, our analysis focused on the most disaggregated level of the data (i.e., SKUs), where the potential advantages of retraining less frequently are likely to be most pronounced. To ensure consistency with the evaluation setup described in Section 3.3, we applied a filtering criterion: for the daily data (M5), only time series with at least two years of observations (730 data points) were retained; for the weekly data (VN1), we considered only those series with at least three years of data (157 observations).
@@ 3.2 Forecasting models
In our study, we focused exclusively on analyzing the stability of global methods, as our primary goal was to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining 111 The analysis of the stability of local models was out of the scope of our study. . Cross-learning has become the standard in many industries dealing with large-scale time series data, particularly in retail demand forecasting, where regular forecasts for thousands of SKUs are required (Januschowski et al., 2020). Therefore, to conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures. The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
@@
Machine learning models have proven effective in forecasting tasks due to their ability to capture complex non-linear relationships in data. They are also relatively easy to train and tend to deliver strong performance when leveraging cross-learning techniques. In this study, we experimented with Linear (Pooled) Regression and four widely used tree-based methods. While machine learning models are easier to train compared to deep learning alternatives, they often rely heavily on high-quality feature engineering (Januschowski et al., 2022). For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M5 and VN1 solutions. We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week). We also included static metadata such as store, product, category, and location identifiers. For the M5 dataset, additional external features like special events were also incorporated. Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective libraries. We compared five machine learning models for time series forecasting. Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso & Hyndman, 2021), (Godahewa et al., 2021b)). Random Forest (RF), an ensemble learning method, captures non-linear patterns and is robust to overfitting, making it suitable for demand forecasting ((Breiman, 2001), (Januschowski et al., 2022)). Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen & Guestrin, 2016), (Ke et al., 2017), Makridakis et al. (2022a)). Categorical Boosting (CatBoost) specializes in handling categorical data with minimal preprocessing (Prokhorenkova et al., 2018).
@@
Deep learning models have gained significant traction in time series forecasting due to their ability to model long-range temporal dependencies and to learn hierarchical representations directly from raw input data (Goodfellow et al., 2016). Unlike machine learning models, deep learning architectures typically do not require extensive manual feature engineering. They can autonomously learn lagged structures, trend, and seasonality directly from the raw series. Nevertheless, training deep models is often more complex due to the higher number of hyperparameters and their sensitivity to configuration choices, which can significantly affect performance (Smyl, 2020). In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates, together with historical values, as inputs. Top solutions’ guidelines were adopted for setting the model hyperparameters when available, and otherwise, the recommended defaults from the respective libraries were used. We compared five deep learning models for time series forecasting. Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter & Schmidhuber, 1997)). Temporal Convolutional Networks (TCN) offer an alternative by using causal convolutions to capture long-range dependencies (Van den Oord et al., 2016). Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
@@
Ensemble learning combines the predictions of multiple base models to improve forecast accuracy, particularly when individual models capture complementary patterns in the data. In the context of global forecasting, ensembles are especially valuable for mitigating the risk of model instability and overfitting, which can occur when relying on a single method (Wang et al., 2023). Following Zanotti (2025a), we estimated four different ensemble models, adopting a simple averaging strategy, which is widely used for its ease of implementation and effectiveness in improving forecast robustness (Claeskens et al., 2016). Indeed, simple averaging can be used for both point and probabilistic forecasting, where equally weighting quantiles has been shown to yield robust results Busetti (2017). To evaluate the potential benefits of ensembling, we simulated a range of ensemble models to assess how forecast stability across the full test window would have differed if combinations had been employed in place of individual base models. Specifically, we constructed simple mean ensembles by aggregating the forecasts of subsets of global models, selecting only those base learners that achieved the highest individual accuracy. This design leverages the strengths of top-performing models while aiming to mitigate their individual weaknesses through averaging. Ensembles of increasing size were considered, combining the top 2, 3, 4, and 5 models according to each out-of-sample performance. Such selective pooling is essentially a variant of model pooling in the sense of Kourentzes et al. (2019), where only a subset of candidate models contributes to the final forecast. Nevertheless, to maintain a balance between potential improvements and the diminishing returns or additional complexity typically associated with larger ensembles (Wang et al., 2023), we limited the maximum ensemble size to five models (out of the eight or ten available, depending on the dataset). The ensemble models were finally validated using the same rolling origin evaluation framework described in Section 3.3. By incorporating ensemble learning into our analysis, we aimed to study whether combining global models could further enhance forecasting stability under different retraining scenarios.